OpenStudy (anonymous):

Suppose $1,000 is divided into two bank accounts. One account pays 1% simple interest per year and the other pays 2.5%. After three years there is a total of $55 in interest between the two accounts. How much was invested into the bank account that pays 2.5% simple interest (rounded to the nearest cent)? $333.33 $444.44 $555.56 $451.55 @iGreen

OpenStudy (igreen):

Okay. Account w/ 1% = x After 3 years: \(x(0.01)(3) \rightarrow 0.03x\) 2.5% account = 1000 - x Earning: (1000 - x)(0.25)(3) = 75 - 0.75x Total interest = 55 That means: 0.03x + 75 -0.075x = 55 Solve for x: x = 444.44 Now subtract 1000 - 444.44 That will be your answer Can you do that? @DUDE..LOOK..ITS..SGTC

OpenStudy (anonymous):

thanks One number is two more than a second number. Twice the first number is triple the second number minus 4. Find the numbers. (negative-2, 0) (0, 2) (10, 8) (5, 7)

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@gorv

OpenStudy (igreen):

x = First Number y = Second Number x = 2 + y 2x = 3y - 4 You have two equations..you can substitute 2 + y for x in the second equation: 2x = 3y - 4 2(2 + y) = 3y - 4 Distribute the 2: 4 + 2y = 3y - 4 Subtract 4 to both sides: 2y = 3y - 8 Subtract 3y to both sides: -y = -8 Multiply -1 to all terms: y = 8 So our second number is 8. Now we can plug in y = 8 into any of the two equations: x = 2 + y x = 2 + 8 Add: x = 10 So what's our answer? @DUDE..LOOK..ITS..SGTC

OpenStudy (anonymous):

(10,8)

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

The length of a rectangle is three more than twice its width. The perimeter is 36 inches. Find the length and width. w = 5, l = 13 w = 8, l = 10 w = 6, l = 12 w = 7, l = 11 @iGreen

OpenStudy (igreen):

l = 2w + 3 Perimeter = 36. Formula for Perimeter of triangle: \(P = 2l + 2w\) Plug in the Perimeter and 2w + 3 for l: \(36 = 2(2w + 3) + 2w\) Distribute: \(36 = 4w + 6 + 2w\) Add: \(36 = 6w + 6\) Subtract 6 to both sides: \(30 = 6w\) Divide: \(w = 5\) So the width is 5. We can plug this back into the equation with the Perimeter: \(P = 2l + 2w\) \(36 = 2l + 2(5)\) Distribute: \(36 = 2l + 10\) Subtract 10 to both sides: \(26 = 2l\) Divide: \(l = 13\) So what's our final answer? @DUDE..LOOK..ITS..SGTC

OpenStudy (anonymous):

(5,13)

OpenStudy (igreen):

Yep. Width = 5 Length = 13